OpenStudy (anonymous):

An archaeologist in Turkey discovers a spear head that contains 32% of it's original amount of C-14. Find the age of the spear head to the nearest year

OpenStudy (anonymous):

Assuming that the amount of C-14 decreases by a rate constant on 1... \[x = N \times e ^{-t}\] x = current amount of C-14 N = original amount of C-14 t = time in years It is stated that x = 0.32 x N Therefore, \[0.32 \times N = N \times e ^{-t}\] Divide both sides by N, \[0.32 = e ^{-t}\] using natural logarithms, \[\ln 0.32 = \ln e ^{-t}\] \[\ln 0.32 = -t \times \ln e, where \ln e=1\] So t = - ln 0.32 = 1.139.. = 1 (to the nearest year) I'm not sure if the answer is correct because i have assumed that t is already in years and that the rate constant of this exponential decay is 1.